OpenStudy (anonymous):

2tan^2xsinx-tan^2x=0

OpenStudy (anonymous):

2tan^2xsinx=tan^2x 2sinx=tan^2x/tan^2x 2sinx=1 sinx=1/2 x=pi/6, 5pi/6

OpenStudy (anonymous):

Just for clarification when you move the tangents to one side they cancel out and all that's left to solve is the sin part right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Wait so where does the two go??

OpenStudy (anonymous):

Nvr I see where it went

OpenStudy (freckles):

\[2 \tan^2(x) \sin(x)-\tan^2(x)=0 \\ \text{ factor left hand expression } \\ \tan^2(x)[2 \sin(x)-1]=0 \\ \text{ two equations to solve } \tan^2(x)=0 \text{ or } 2\sin(x)-1=0\]